1. Find integers s, t such that 172s + 151t == 1. For full credit, find the smallest magnitude integers you can. Hint: First find gcd(151, 172) = 1.
Answer #1

`Hey,

Note: Brother in case of any queries, just comment in box I would be very happy to assist all your queries

By using extended eulidean algorithm,

Set up a division problem where a is larger than b.
a ÷ b = c with remainder R. Do the division. Then replace a with b, replace b with R and repeat the division. Continue the process until R = 0.

172 ÷ 151 = 1 R 21    (172 = 1 × 151 + 21)

151 ÷ 21 = 7 R 4    (151 = 7 × 21 + 4)

21 ÷ 4 = 5 R 1    (21 = 5 × 4 + 1)

4 ÷ 1 = 4 R 0    (4 = 4 × 1 + 0)

So, s=36, t=-41

Kindly revert for any queries

Thanks.